RC Update and Run Plan for Aug 14, 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, remote

Wednesday Aug 14 2020.

General updates

  • Hydrogen and empty Reference cell runs were taken during SWING shift (08/13) [1]
  • Polarized He3 data taken during OWL shift. We optimized the current/prescales for SHSM and HMS
  • Works out to 24 mA---- prescale 2 for SHMS and prescale 1 for HMS
  • Beam current / prescale values for both HMS and SHMS were optimized to keep them around 4 kHZ (about ~85% live time for SHMS).
  • Were down for half of the day (10:30 -14:30 ): https://cebaf.jlab.org/dtm/open-events?event_id=6951
  • Melanie worked on optimizing the QWP: Transverse Left QWP
  • Brad replaced the UPS, hopefully fixed the DC gas supply pressure alarm Gas_Supply_Alarm
  • Continuing to coordinate with Hall A about IHWP. They will nominally change every 8 hours. Will remind the shift crew to notify me if it hasn't occurred in 8 hours.
  • 3:30 - Jay called to ask if we are interested in switching to 1st pass beam today. Hall A's Moller quad power supplies need replaced and will be down for some time.
  • We will keep an eye on how long they will be down for and may make the pass change on Monday or Tuesday (08/17 or 08/18) of next week.....

Pending problem

  • Beam trip rate

Recent polarization measurements

  • 43% at 3:30 (no fit error but Bill will look into it on SWING)
  • Fitting Issue: Estimate around 40% at 10:00 DAY_NMR
  • 40.8% at 05:00 OWL_NMR
  • 41.7% at 00:05 SWING_NMR

Current Run Plan

Normal Production running

  • Prerequisites
    • Make sure the raster size is correct (4-4.5)
    • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
    • Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
    • Always take 120Hz of EDTM
    • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
    • Change the IHWP at least every 8 hours when running on glass cells. If Hall A hasn't changed in 8 hours, call the RC
  • Run plan
    • Start back with the Beam / Raster checkout (short procedure) Checkout
    • Production on 3He; if not able to reach requested current, pro-rate the ABUs

Notes for tomorrow's 7:45 meeting

  • Beam stability / trip rate?

Future

  • We are on the schedule to switch to 1st pass on Monday (Aug 24) and switch back to 5th pass on Thursday (Aug 27).
  • Check the SAD calender for schedule updates (MCC white board isn't currently being updated)
  • Next maintenance is hasn't been scheduled as of yet
  • EPR diode good for now, but check for future planned accesses.
  • If Hall will have opening for 1+ hours, let Simona know as there is work to do on the cerenkov.

Hall Access

When there is a down:

Notices in Effect for Shift Crew

  • Beam positions: [2]
  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[3]



Planned Accelerator Configuration Changes

  • Hall C to 1st pass on 8/24 and back to 5th pass on 8/27


  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920